OpenStudy (anonymous):

Compute the derivative function f '(x) algebraically. Explain how please? f(x) = x^2 − 6

OpenStudy (anonymous):

f'(x) = 2x

OpenStudy (anonymous):

u use the formula if f(x) = ax^n f'(x) = anx^ (n-1)

OpenStudy (anonymous):

I dont think that here this is what was asked. You should use the definition of the derivative f'(x)=lim (f(x+h)-f(x))/h as h tends to 0 In this case lim (x+h)^2-6-x^2+6/h=lim(2xh+h^2)/h=lim 2x +h= 2x (as h tends to 0)

OpenStudy (amistre64):

first principles is what i believe its asking for too